# Combining Evidence Across Filtrations

### v3

@@ 1.1 Background: E-processes and sequential anytime-valid inference (SAVI)
[…]
@@
We provide a glossary of the key definitions and notations in Section A.
@@ 1.2 A motivating example from financial time series modeling
$\mathcal{H}_{0}^{\mathsf{IID}}:X_{1},X_{2},\dotsc\;\text{are IID.}$. (2)
@@
[…]
The manager now faces an intriguing challenge. By design, the UI e-process is powerful against Markovian alternatives but powerless against changepoint alternatives; the opposite is true for the conformal test martingale. Yet, unlike in the case of e-values, these two e-processes cannot be averaged straightforwardly because the conformal test martingale only has restricted anytime-validity, only within a sub-filtration of the data filtration.
@@
For each fixed $t$, $\mathbb{E}_{P}[\bar{m}_{t}]\leq 1$ under any $P\in\mathcal{H}_{0}^{\mathsf{IID}}$, so $\bar{m}$ is a sequence of e-values. However, is $\bar{m}$ an e-process? To answer this question, we need to clarify which filtration we are working with. The most natural candidate is the data filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$, where $\mathcal{F}_{t}=\sigma(X_{1},\dotsc,X_{t})$. However, while the UI e-process $\mathfrak{e}^{\mathsf{UI}}$ is a well-defined e-process in $\mathbb{F}$, the conformal test martingale $\mathfrak{e}^{\mathsf{conf}}$ is not an e-process in $\mathbb{F}$. In fact, $\mathfrak{e}^{\mathsf{conf}}$ is only an e-process in a sub-filtration $\mathbb{G}=(\mathcal{G}_{t})_{t\geq 0}\subsetneq\mathbb{F}$ constructed by the conformal p-values—that is, $\mathcal{G}_{t}=\sigma(s_{1},\dotsc,s_{t})$. In particular, $\mathfrak{e}^{\mathsf{conf}}$ is only valid at arbitrary $\mathbb{G}$-stopping times but not at arbitrary $\mathbb{F}$-stopping times. Therefore, $\bar{m}$ is a sequence of e-values that is not an e-process in $\mathbb{F}$. In fact, even in the sub-filtration $\mathbb{G}$, $\bar{m}$ is not an e-process because its first term, $\mathfrak{e}_{t}$, is not $\mathcal{G}_{t}$-measurable.
@@
While $\tau^{\mathbb{F}}$ is a $\mathbb{F}$-stopping time, it is not a $\mathbb{G}$-stopping time because it depends directly on the individual data points and not just the conformal p-values $s_{1},\dotsc,s_{t}$. Over 100 repeated runs, we compute both the UI e-process and the conformal test martingale until they each reach $\tau^{\mathbb{F}}$. Figure 1 shows that, while $\mathfrak{e}^{\mathsf{UI}}_{\tau^{\mathbb{F}}}$ remains small ($\leq 1$) across repeated runs, $\mathfrak{e}^{\mathsf{conf}}_{\tau^{\mathbb{F}}}$ often exceeds $1$. Over $10,000$ repeated runs, we find that the mean of $\mathfrak{e}_{\tau^{\mathbb{F}}}^{\mathsf{conf}}$ is significantly larger than one ($1.33\pm 0.016$), while the mean of $\mathfrak{e}_{\tau^{\mathbb{F}}}^{\mathsf{UI}}$ remains small ($0.25\pm 0.008$). This example makes clear that an e-process in a coarser filtration ($\mathfrak{e}^{\prime}$ in $\mathbb{G}$) need not be an e-process in a finer filtration ($\mathbb{F}$).
@@
Figure 1: E-processes in a sub-filtration $\mathbb{G}\subsetneq\mathbb{F}$ are not valid at $\mathbb{F}$-stopping times. The plot shows 100 instances the UI e-process (solid blue) and the conformal test martingale (dashed orange) for testing randomness ($\mathcal{H}_{0}^{\mathsf{IID}}$), using data generated repeatedly from an IID $\mathsf{Ber}(0.3)$. Each e-process is stopped at an $\mathbb{F}$-stopping time $\tau^{\mathbb{F}}$, the first time we observe five consecutive zeros in the data (4). For the conformal test martingale, many stopped e-values are above $1$, even though the data comes from a null distribution; in contrast, the UI e-process remains small ($\leq 1$) at $\tau^{\mathbb{F}}$. Over $10,000$ repeated runs, the mean of stopped e-values for the conformal test martingale (vertical dashed orange) is $1.33\pm 0.016$, implying that $\mathbb{E}[\mathfrak{e}_{\tau^{\mathbb{F}}}^{\mathsf{conf}}]>1$.
@@
This phenomenon poses both a technical and practical challenge when combining sequentially obtained evidence across different filtrations. On the technical side, this puts an emphasis on specifying an e-process’ underlying filtration, which is often “swept under the rug.” On the practical side, the experimenter must be cautious about exactly what information is included in the stopping rule because accessing extra information can (either inadvertently or deliberately) invalidate the resulting test, particularly when that information “leaks” into the decision to stop the experiment. This is impractical in a setup like Example 1, as the experimenter then cannot look at the data itself.
@@
Some recent works have noted the challenge of combining e-processes (or test martingales) across filtrations. For randomness testing, Vovk et al., 2022a (, section 9.3 in 2nd ed.) note that the mean of conformal test martingales need not be a test martingale, even when computed on the same data, as each one is constructed using differently randomized conformal p-values that yield different filtrations. Besides randomness testing, there are at least two other practical examples in the literature, one in sequential forecast comparison (Henzi and Ziegel,, 2022; Choe and Ramdas,, 2023) and another in independence testing (Podkopaev et al.,, 2023; Henzi and Law,, 2024); we discuss these examples in Section 5. Yet another example is testing a scale-invariant Gaussian mean, for which different e-processes exist in different filtrations (Pérez-Ortiz et al.,, 2024; Wang and Ramdas, 2025a, ). We review this example in Section C.
@@
Even though we focused on practical motivation thus far, our main results have standalone implications in enhancing our theoretical understanding of both adjusters and e-processes. First, we establish a novel characterization of adjusters that suggests adjusters are not only sufficient but also, in a formal sense, necessary for lifting arbitrary e-processes. Second, our “$\mathfrak{e}$-lifting” result implies that, unlike in the case of test martingales, there always exists a powerful e-process for a testing problem in the data filtration whenever there exists one in a coarser filtration.
@@ 1.3 Related work on combining e-values (and p-values)
The rest of the paper is organized as follows. In Section 2, we briefly review other relevant notions from anytime-valid inference, including p-processes and sequential tests, and then formally discuss both calibrators and adjusters. In Section 3, we introduce the $\mathfrak{p}$-lifting and $\mathfrak{e}$-lifting procedures and prove that the lifted p-/e-processes are valid at stopping times in the finer filtration. In Section 4, we describe experimental results on combining arbitrary e-processes for testing randomness, using both simulated data and real-world financial time series data. In Section 5, we describe two other applications of $\mathfrak{e}$-lifting, one in independence testing and another in sequential forecast comparison. Then, in Section 6, we present our main characterization theorem for adjusters that formalizes a sense in which using adjusters is necessary for lifting arbitrary e-processes across filtrations. Finally, in Section 7, we discuss additional theoretical and practical implications of our main results.
@@ 2.1 SAVI part 2: P-processes, sequential tests, and Ville’s inequality
[…]
@@
[…]
@@ 2.2 Calibrating e-processes and p-processes
Combining evidence in the forms of an e-process and a p-process requires calibrating one to another. Prior work by Shafer et al., (2011); Vovk and Wang, (2021) suggests that any e-process can be calibrated into a p-process and vice versa. Hereafter, we allow e-processes and calibrators to be defined on the extended nonnegative reals, $[0,\infty]$, and we refer the reader to Wang and Ramdas, 2025b (, section 2.2) for the arithmetic and measure-theoretic properties on $[0,\infty]$.
@@
Converting a p-process into an e-process requires using a p-to-e calibrator (Vovk,, 1993; Sellke et al.,, 2001; Shafer et al.,, 2011; Ramdas et al.,, 2020; Vovk and Wang,, 2021). Formally, a decreasing function $\mathsf{C}:[0,1]\to[0,\infty]$ is a p-to-e calibrator if
[…]
@@
[…]
@@ 2.3 Adjusting the running maximum of e-processes
In general, the running maximum of a test supermartingale or an e-process is not a test supermartingale or an e-process; in fact, the expected running maximum will typically diverge to infinity under the null (Ramdas et al.,, 2022). For test supermartingales, Dawid et al., 2011a ; Dawid et al., 2011b ; Shafer et al., (2011) develop a class of functions called adjusters, which convert the running maximum of a test supermartingale into an e-process (in the same filtration). The papers use the terms simple lookback adjusters, capital calibrators, and martingale calibrators, respectively; here, we call them adjusters for brevity. In the context of statistical inference, adjusters allow us to “insure against the loss of evidence” by trading off some amount of evidence. This is useful when, e.g., the test supermartingale sharply decreases following a changepoint.
@@
[…]
Furthermore, $\mathsf{A}$ is admissible (i.e., no other adjuster dominates it on the entire domain of $[1,\infty]$) if and only if it is right-continuous, $\mathsf{A}(\infty)=\infty$, and the condition (11) holds with equality.
@@
[…]
@@
In Figure 2, we plot four admissible adjusters (in logarithmic scale): $\mathsf{A}_{\mathsf{mix}}$, $\mathsf{A}_{\mathsf{KV}}$, $\mathsf{A}_{\mathsf{sqrt}}$, and $\mathsf{A}_{\kappa}$ for $\kappa=1/2$. While all of these adjusters are admissible, it is apparent some grow substantially faster than others: $\mathsf{A}_{\mathsf{mix}}$ and $\mathsf{A}_{\mathsf{KV}}$ grow linearly in $\mathfrak{e}$, up to logarithmic factors, and thus they eventually dominate the other three variants, which only grow sublinearly in $\mathfrak{e}$. $\mathsf{A}_{\mathsf{mix}}$ and $\mathsf{A}_{\mathsf{KV}}$ are close to each other, although $\mathsf{A}_{\mathsf{mix}}$ dominates $\mathsf{A}_{\mathsf{KV}}$ for $\mathfrak{e}\geq 4.40$. Thus, we use the mixture adjuster $\mathsf{A}_{\mathsf{mix}}$ as a default. In Section D.1, we discuss another variant due to Shafer et al., (2011) that grows faster than $\mathsf{A}_{\mathsf{mix}}$ by a constant factor by completely zeroing out small values of $\mathfrak{e}$.

### v4

@@ 1.1 Background: E-processes and sequential anytime-valid inference (SAVI)
$\text{for any $\mathbb{F}$-stopping time $\tau$ and any $P\in\mathcal{P}$,}\quad\mathbb{E}_{P}[\mathfrak{e}_{\tau}]\leq 1.$ (1)
@@
We provide a glossary of the key definitions and notations in Section A. For a more comprehensive overview, we refer the reader to section 7 of Ramdas and Wang, (2025).
@@ 1.2 A motivating example from financial time series modeling
$\mathcal{H}_{0}^{\mathsf{IID}}:X_{1},X_{2},\dotsc\;\text{are IID.}$ (2)
@@
A priori, the manager does not know exactly how the data may deviate from randomness. Thus, they will want to leverage (at least) two different testing procedures that achieve power under different alternatives. One such procedure is based on the universal inference (UI; Wasserman et al.,, 2020; Ramdas et al.,, 2022) e-process, denoted as $\mathfrak{e}^{\mathsf{UI}}=(\mathfrak{e}_{t}^{\mathsf{UI}})_{t\geq 0}$, which generalizes likelihood ratios to composite nulls and alternatives. Formally, the UI e-process can be written as $\mathfrak{e}_{t}^{\mathsf{UI}}=\frac{\int_{\theta\in[0,1]^{2}}Q_{\theta}(X_{1},\dotsc,X_{t})p(\theta)d\theta}{\sup_{\mu\in[0,1]}P_{\mu}(X_{1},\dotsc,X_{t})}$, where $P_{\mu}$ is the IID $\mathsf{Ber}(\mu)$ distribution; $Q_{\theta}$ is a first-order Markov distribution with transition parameters $\theta=(p_{0\to 1},p_{1\to 1})\in[0,1]^{2}$; and $p(\theta)$ is a “prior” distribution over $\theta$. Another is based on the conformal test martingale (CTM) (Vovk, 2021b, ), denoted as $\mathfrak{e}^{\mathsf{conf}}=(\mathfrak{e}_{t}^{\mathsf{conf}})_{t\geq 0}$, where $\mathfrak{e}_{t}^{\mathsf{conf}}=\prod_{i=1}^{t}[1+\lambda(s_{i}-1/2)]$ for $\lambda\in\mathbb{R}$ and “conformal p-values” $s_{1},\dotsc,s_{t}\in[0,1]$. Intuitively, these conformal p-values capture pointwise deviations from the running mean of the data. Under the null, they are IID $\mathsf{Unif}[0,1]$ with mean $1/2$; under a changepoint alternative, their mean deviates from $1/2$ (to 0 or 1). Further details about these e-processes are not central to our discussion, so we defer them to Section B.
The manager now faces an intriguing challenge. By design, the UI e-process is powerful against Markovian alternatives but powerless against changepoint alternatives; the opposite is true for the CTM. Yet, unlike in the case of e-values, these two e-processes cannot be averaged straightforwardly because the CTM only has restricted anytime-validity, only within a sub-filtration of the data filtration.
@@
For each fixed $t$, $\mathbb{E}_{P}[\bar{m}_{t}]\leq 1$ under any $P\in\mathcal{H}_{0}^{\mathsf{IID}}$, so $\bar{m}$ is a sequence of e-values. However, is $\bar{m}$ an e-process? To answer this question, we need to clarify which filtration we are working with. The most natural candidate is the data filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$, where $\mathcal{F}_{t}=\sigma(X_{1},\dotsc,X_{t})$. However, while the UI e-process $\mathfrak{e}^{\mathsf{UI}}$ is an e-process in $\mathbb{F}$, the conformal test martingale $\mathfrak{e}^{\mathsf{conf}}$ is not. In fact, $\mathfrak{e}^{\mathsf{conf}}$ is only an e-process in a sub-filtration $\mathbb{G}=(\mathcal{G}_{t})_{t\geq 0}\subsetneq\mathbb{F}$ constructed by the conformal p-values—that is, $\mathcal{G}_{t}=\sigma(s_{1},\dotsc,s_{t})$. Note that $\mathbb{G}$ is a strictly coarser filtration than $\mathbb{F}$—for example, neither the sequence of events $\{\sum_{i=1}^{t}X_{i}\geq c\}_{t\geq 0}$ nor the raw data sequence $(X_{t})_{t\geq 0}$ is adapted to $\mathbb{G}$. Thus, $\mathfrak{e}^{\mathsf{conf}}$ is only valid at arbitrary $\mathbb{G}$-stopping times but not at arbitrary $\mathbb{F}$-stopping times. This means that $\bar{m}$ is a sequence of e-values that is not an e-process in $\mathbb{F}$; even in $\mathbb{G}$, $\bar{m}$ is not an e-process because its first term, $\mathfrak{e}_{t}^{\mathsf{UI}}$, is not $\mathcal{G}_{t}$-measurable.
@@
While $\tau^{\mathbb{F}}$ is a $\mathbb{F}$-stopping time, it is not a $\mathbb{G}$-stopping time because it depends directly on the individual data points and not just the conformal p-values $s_{1},\dotsc,s_{t}$. In Figure 1, over 100 repeated runs, we plot both the UI e-process and the CTM until they each reach $\tau^{\mathbb{F}}$. While $\mathfrak{e}^{\mathsf{UI}}_{\tau^{\mathbb{F}}}$ remains small ($\leq 1$) across repeated runs, $\mathfrak{e}^{\mathsf{conf}}_{\tau^{\mathbb{F}}}$ often exceeds $1$. Over $10,000$ repeated runs, we find that the mean of $\mathfrak{e}_{\tau^{\mathbb{F}}}^{\mathsf{conf}}$ is significantly larger than one ($1.39\pm 0.017$), while the mean of $\mathfrak{e}_{\tau^{\mathbb{F}}}^{\mathsf{UI}}$ remains small ($0.25\pm 0.008$). This example makes clear that an e-process in a coarser filtration ($\mathfrak{e}^{\prime}$ in $\mathbb{G}$) need not be an e-process in a finer filtration ($\mathbb{F}$).
@@
Figure 1: E-processes in a sub-filtration $\mathbb{G}\subsetneq\mathbb{F}$ are not valid at $\mathbb{F}$-stopping times. The plot shows 100 instances of the UI e-process $\mathfrak{e}^{\mathsf{UI}}$ (solid blue) and the conformal test martingale $\mathfrak{e}^{\mathsf{conf}}$ (dashed orange; $\lambda=1$) for testing randomness ($\mathcal{H}_{0}^{\mathsf{IID}}$), using data generated repeatedly from an IID $\mathsf{Ber}(0.3)$. Each e-process is stopped at an $\mathbb{F}$-stopping time $\tau^{\mathbb{F}}$, the first time we observe five consecutive zeros in the data (4). For the conformal test martingale, many stopped e-values are above 11, even though the data comes from a null distribution; in contrast, the UI e-process remains small ($\leq 1$) at $\tau^{\mathbb{F}}$. Over $10,000$ repeated runs, the mean of stopped e-values for the conformal test martingale (vertical dashed orange) is $1.39\pm 0.017$, implying that $\mathbb{E}[\mathfrak{e}_{\tau^{\mathbb{F}}}^{\mathsf{conf}}]>1$.
@@
This phenomenon poses both a technical and practical challenge when combining sequentially obtained evidence across different filtrations. On the technical side, this puts an emphasis on specifying an e-process’ underlying filtration, which is often “swept under the rug.” On the practical side, the experimenter must be cautious about exactly what information is included in the stopping rule because accessing extra information can (either inadvertently or deliberately) invalidate the resulting test, particularly when that information “leaks” into the decision to stop the experiment. This is impractical, as the experimenter cannot then look at the data itself.
At this point, we note that the broader takeaway remains the same even if we substituted the CTM with its improved variants in the literature, such as the “simple jumper” CTM (e.g., Vovk et al., 2022a, ) that we describe in Section B.3. Other variants of the CTM are still invalid under $\mathbb{F}$-stopping times, and it is known that CTMs that are powerful against changepoint alternatives do not achieve power against Markovian alternatives (fig. 9.8 of Vovk et al., 2022a, ). An analogous remark could be made about the UI e-process (e.g., Saha and Ramdas,, 2024).
@@
Some recent works have noted the challenge of combining e-processes (or test martingales) across filtrations. For randomness testing, Vovk et al., 2022a (, section 9.3 in 2nd ed.) note that even the average of the same CTMs across different runs need not be a test martingale, as each one is constructed using differently randomized conformal p-values that yield different filtrations. Besides randomness testing, there are at least two other practical examples in the literature, one in sequential forecast comparison (Henzi and Ziegel,, 2022; Choe and Ramdas,, 2023) and another in independence testing (Podkopaev et al.,, 2023; Henzi and Law,, 2024); we discuss these examples in Section 5. Yet another example is testing a scale-invariant Gaussian mean, for which different e-processes exist in different filtrations (Pérez-Ortiz et al.,, 2024; Wang and Ramdas, 2025a, ). We review this example in Section C.
@@
Even though we focused on practical motivation thus far, our main results have standalone implications in enhancing our theoretical understanding of both adjusters and e-processes. First, we establish a novel characterization of adjusters that suggests adjusters are not only sufficient but also, in a formal sense, necessary for lifting arbitrary e-processes. Second, our $\mathfrak{e}$-lifting” result implies that, unlike in the case of test martingales, there always exists a powerful e-process for a testing problem in the data filtration whenever there exists one in a coarser filtration.
@@ 1.3 Related work on combining e-values (and p-values)
The rest of the paper is organized as follows. In Section 2, we briefly review other relevant notions in SAVI, including p-processes and sequential tests, and then formally discuss calibrators and adjusters. In Section 3, we introduce the $\mathfrak{p}$-lifting and $\mathfrak{e}$-lifting procedures and prove that the lifted p-/e-processes are valid at stopping times in finer filtrations. In Section 4, we describe experimental results on combining arbitrary e-processes for testing randomness, using both simulated data and real-world financial time series data. In Section 5, we describe two other applications of $\mathfrak{e}$-lifting, one in independence testing and another in forecast comparison. In Section 6, we present a characterization theorem for adjusters that formalizes a sense in which using adjusters is necessary for lifting arbitrary e-processes across filtrations. Finally, in Section 7, we discuss theoretical and practical implications of our main results.
@@ 2.1 SAVI part 2: P-processes, sequential tests, and Ville’s inequality
$\text{for any $\mathbb{F}$-stopping time $\tau$ and any $P\in\mathcal{P}$,}\quad P(\mathfrak{p}_{\tau}\leq\alpha)\leq\alpha.$ (5)
@@
$\text{for any $\mathbb{F}$-stopping time $\tau$ and any $P\in\mathcal{P}$,}\quad P(\phi_{\tau}=1)\leq\alpha.$ (6)
@@ 2.2 Calibrating e-processes and p-processes
Combining evidence in the forms of an e-process and a p-process requires calibrating one to another. Prior work by Shafer et al., (2011); Vovk and Wang, (2021) suggests that any e-process can be calibrated into a p-process and vice versa. Hereafter, we allow e-processes and calibrators to be defined on the extended nonnegative reals, $0,\infty]$, and we refer the reader to Wang and Ramdas, 2025b (, section 2.2) for the arithmetic and measure-theoretic properties on $0,\infty]$.
@@
In the opposite direction, p-to-e calibrators (Vovk,, 1993; Sellke et al.,, 2001; Shafer et al.,, 2011; Ramdas et al.,, 2020; Vovk and Wang,, 2021) allow us to convert a p-process into an e-process. A decreasing function $\mathsf{C}:[0,1]\to[0,\infty]$ is a p-to-e calibrator if
$\int_{0}^{1}\mathsf{C}(\mathfrak{p})d\mathfrak{p}\leq 1,$ (9)
@@
$\mathsf{C}_{\mathsf{mix}}(\mathfrak{p})=\int_{0}^{1}\mathsf{C}_{\kappa}(\mathfrak{p})d\kappa=\frac{1-\mathfrak{p}+\mathfrak{p}\log(\mathfrak{p})}{\mathfrak{p}\log^{2}(\mathfrak{p})},\quad\forall\mathfrak{p}\in[0,1).$ (10)
@@ 2.3 Adjusting the running maximum of e-processes
In general, the running maximum of a test martingale or an e-process is not an e-process (Ramdas et al.,, 2022); in fact, the running maximum of a test martingale, in expectation under the null, can tend to infinity (we describe a concrete example in Section D). For test (super)martingales, Dawid et al., 2011a ; Dawid et al., 2011b ; Shafer et al., (2011) develop a class of functions called adjusters, which convert the running maximum of a test (super)martingale into an e-process. The papers use the terms simple lookback adjusters, capital calibrators, and martingale calibrators, respectively; here, we call them adjusters for brevity. In the context of statistical inference, adjusters allow “insuring against the loss of evidence” by trading off some amount of evidence. This is useful when, e.g., the test martingale sharply decreases following a changepoint.
@@
$\int_{1}^{\infty}\frac{\mathsf{A}(\mathfrak{e})}{\mathfrak{e}^{2}}d\mathfrak{e}\leq 1.$ (11)
Furthermore, $\mathsf{A}$ is admissible (i.e., no other adjuster dominates it on the entire domain of $1,\infty]$) if and only if it is right-continuous, $\mathsf{A}(\infty)=\infty$, and the condition (11) holds with equality.
@@
$\mathsf{A}_{\mathsf{mix}}(\mathfrak{e})=\int_{0}^{1}\mathsf{A}_{\kappa}(\mathfrak{e})d\kappa=\frac{\mathfrak{e}-1-\log(\mathfrak{e})}{\log^{2}(\mathfrak{e})},\quad\forall\mathfrak{e}>1.$ (12)
@@
In Figure 2, we plot four admissible adjusters (in logarithmic scale): $\mathsf{A}_{\mathsf{mix}}$, $\mathsf{A}_{\mathsf{KV}}$, $\mathsf{A}_{\mathsf{sqrt}}$, and $\mathsf{A}_{\kappa}$ for $\kappa=1/2$. While all of these adjusters are admissible, it is apparent some grow substantially faster than others: $\mathsf{A}_{\mathsf{mix}}$ and $\mathsf{A}_{\mathsf{KV}}$ grow linearly in $\mathfrak{e}$, up to logarithmic factors, and thus they eventually dominate the other three variants, which only grow sublinearly in $\mathfrak{e}$. $\mathsf{A}_{\mathsf{mix}}$ and $\mathsf{A}_{\mathsf{KV}}$ are close to each other, although $\mathsf{A}_{\mathsf{mix}}$ dominates $\mathsf{A}_{\mathsf{KV}}$ for $\mathfrak{e}\geq 4.40$. Thus, we use the mixture adjuster $\mathsf{A}_{\mathsf{mix}}$ as a default. In Section E.1, we discuss another variant due to Shafer et al., (2011) that grows faster than $\mathsf{A}_{\mathsf{mix}}$ by a constant factor by completely zeroing out small values of $\mathfrak{e}$.